OpenStudy (anonymous):

7. Find the area of a parallelogram with base 8 units and altitude 4.5 units.

OpenStudy (anonymous):

area of parallelogram = b * h= 8 * 4.5= 36 squa units

OpenStudy (anonymous):

here b = base and h = altitude